OpenStudy (anonymous):

Express the complex number in trigonometric form. 3

OpenStudy (anonymous):

Answer choices: 3(cos 0° + i sin 0°) 3(cos 90° + i sin 90°) 3(cos 180° + i sin 180°) 3(cos 270° + i sin 270°)

OpenStudy (anonymous):

Do you think you could help, @Hero?

OpenStudy (anonymous):

I was thinking it was either 0 or 180 because it doesn't have the imaginary component...

OpenStudy (anonymous):

\(\cos(0)=1\) and \(\sin(0)=0\) so \[3=3\left(\cos(0)+i\sin(0)\right)\]